OpenStudy (anonymous):

3. Solve 3x - 4 = 11 for x. (Points : 4) 2 3 4 5

OpenStudy (ja1):

ok do you know how to start?

OpenStudy (ja1):

Ok so I will start you off then: 3x - 4 = 11 +4 +4

OpenStudy (ja1):

Are you still here?

OpenStudy (ja1):

Ok then moving on now you get: 3x = 15

OpenStudy (ja1):

Now you need to isolate x and so you would divide 3 on both sides

OpenStudy (ja1):

3x = 15 /3 /3

OpenStudy (ja1):

Now you have x = 5

OpenStudy (ja1):

Do you understand or do you need more help?